OpenStudy (anonymous):

A line passes through the point (–2, 4), and its y-intercept is (0, –6). What is the equation of the line that is perpendicular to the first line and passes through the point (5, –4)?

OpenStudy (phebe):

\(\color{RED}{WELCOME~TO~OPENSTUDY~}\)

OpenStudy (anonymous):

Please just help me lol

OpenStudy (phebe):

LOL The equation of the line, in general form, that passes through the point (1, 1) and has a y-intercept of 2 is y=-x+2, or y=2-x

OpenStudy (phebe):

@Owlcoffee

OpenStudy (owlcoffee):

The y-intercept is but a notable point, but that point has to also belong to the line and the y-axis, which is an useful tool to find equation of lines. So, we will just find the slope of the given point and the y-interception of the first line and hopefully use the point-slope formula in order to find it's equation: \[m_1=\frac{ -6-4 }{ 0-(-2) }\]

OpenStudy (anonymous):